OpenStudy (ineedhelpfast12):

solve the linear system using substitution. 2x-y=23 x-9=-1

OpenStudy (wilsondanielle):

the second equation gives you the value of x (-1+9=8)...

OpenStudy (ineedhelpfast12):

y=-7 x=8

OpenStudy (wilsondanielle):

If you HAVE to do substitution : 2(8) - y = 23 16 - y = 23 23-16 = 7 = -y -y = 7

OpenStudy (ineedhelpfast12):

thanks